# Large Language Model Enhanced Recommender Systems: Taxonomy, Trend, Application and Future

### v2

@@ 4.3.3. User & Item
The other line of work supplements the user and item LLM embeddings as extra features to augment the content-based RS models. For instance, Laser (Lin et al., 2024) utilizes the derived user and item LLM embeddings with an adapter for feature enhancement. As for LARR (Wan et al., 2024), it also adopts the LLM embeddings as extra features, but further highlights the importance of fine-tuning the LLM specialized for RS.
@@ 4.4. Embedding Guidance
Due to the lengthy prompt of user interactions, user LLM embeddings are often noisy. Thus, considering the LLM embeddings as guidance is a feasible way for user-only enhancement. LLM4SBR (Qiao et al., 2024) enhances the short-term and long-term preferences for session-based recommendation models by identifying the representative items via LLM embeddings. By comparison, LLM-ESR (Liu et al., 2024e) retrieves similar users by LLM embeddings to enhance the long-tail users during the training. These two works both aim to search for relevant users or items from a semantic view and enhance the training process. Differently, LLM4MSR (Wang et al., 2024g) generates the user-specified parameters for multi-scenario recommendation models by LLM embeddings accompanied with meta networks. As for User & Item line, the research studies guide both the user and item representation learning. RLMRec (Ren et al., 2024b) kicks off this line, which adds an extra loss to align the collaborative user and item embeddings with corresponding LLM embeddings during training. Based on the insights of RLMRec, DaRec (Yang et al., 2024a) further proposes to disentangle the shared and specific parts for both LLM and collaborative embeddings. Only imposing the alignment on shared parts will lead to better performance.
@@ 5. Trend
(Liu et al., 2024a),(Luo et al., 2024c), (Huang et al., 2024a),(Wang et al., 2024d)
(Liu et al., 2024f),(Sun et al., 2024), (Wang et al., 2024f),(Zhang et al., 2024b)
(Hu et al., 2024),(Cui et al., 2024), (Xu et al., 2024a),(Liu et al., 2024g)
@@
(Qiao et al., 2024),(Zhao et al., 2024b)
Vicuna Y (Wang et al., 2024c)
@@
(Wang et al., 2024e)
To investigate the trend of the LLMERS, we visualize the current works based on the type of LLM and semantics in Figure 6. As mentioned before, the key to LLM enhancement lies in the semantics. At the early stage, many works (Xi et al., 2024a; Zhao et al., 2024c) prompt the LLM to derive the natural language, which contains knowledge and semantics for augmentation. Since they are readable and understandable, we denote them as explicit semantics. By comparison, the cluster of implicit semantics means that the mediation from LLM for enhancement is non-verbal, e.g., the hidden states of LLM. Though explicit knowledge should be preferable by humans due to its explainability, we find that more recent works have embraced implicit one. The reason lies in the better performance of implicit knowledge, while the information loss occurs in the extra encoding process for explicit knowledge.
@@
  • Model enhancement has gained more attention recently, because it can combine with implicit semantics and fine-tuned LLM better.
6. Applications and Resources
Traditional recommender systems have been widely adopted in various applications, such as news recommendations. Though LLMERS is also based on traditional RS, it often requires abundant side information, especially informative texts, to make full use of the reasoning and understanding abilities of LLM. Thus, current LLMERS works often focus on applications that have one of two following characteristics: (i) it has a number of normal features that LLM can understand; (ii) it has a volume of texts that LLM can summarize. The E-commerce application belongs to the former one, where users have many profile features and commodities have attributes (Wang et al., 2024a; Liu et al., 2024e). Since the LLM owns open-world knowledge, these features can be understood by it and derive semantics for enhancement. By comparison, the news recommendation is the typical application that utilizes the LLM to summarize the abundant news texts (Liu et al., 2024a). To show it more clearly, we list the typical applications that LLMERS has targeted in Table 2.
Datasets and benchmarks are often the most vital resources for one research field. For the dataset, many have included either informative features or abundant texts for LLM usage. We list the available public datasets with their access links according to the applications in Table 2 to facilitate further research. As for the benchmark, unfortunately, there has been no one since it is an emerging direction. Nevertheless, many research studies have open-sourced their code, which pushes the advancements of this field. We conclude all the available codes in our survey repository  [^1] https://github.com/liuqidong07/Awesome-LLM-Enhanced-Recommender-Systems.

### v3

@@ 4.3.3. User & Item
The other line of work supplements the user and item LLM embeddings as extra features to augment the content-based RS models. For instance, Laser (Lin et al., 2024) utilizes the derived user and item LLM embeddings with an adapter for feature enhancement. LARR (Wan et al., 2024) adopts the LLM embeddings as extra features but further highlights the importance of fine-tuning the LLM specialized for RS.
@@ 4.4. Embedding Guidance
Due to the lengthy prompt of user interactions, user LLM embeddings are often noisy. Thus, considering the LLM embeddings as guidance is a feasible way for user-only enhancement. LLM4SBR (Qiao et al., 2024b) enhances the short-term and long-term preferences for session-based recommendation models by identifying the representative items via LLM embeddings. By comparison, LLM-ESR (Liu et al., 2024e) retrieves similar users by LLM embeddings to enhance the long-tail users during the training. These two works both aim to search for relevant users or items from a semantic view and enhance the training process. Differently, LLM4MSR (Wang et al., 2024i) generates the user-specified parameters for multi-scenario recommendation models by LLM embeddings accompanied by meta networks. As for User & Item line, the research studies guide both the user and item representation learning. RLMRec (Ren et al., 2024b) kicks off this line, which adds an extra loss to align the collaborative user and item embeddings with corresponding LLM embeddings during training. Based on RLMRec’s insights, DaRec (Yang et al., 2024a) proposes to disentangle the shared and specific parts for both LLM and collaborative embeddings. Only imposing the alignment on shared parts will lead to better performance.
@@ 5. Trend
(Liu et al., 2024a),(Luo et al., 2024e), (Huang et al., 2024a),(Wang et al., 2024e)
(Liu et al., 2024f),(Sun et al., 2024), (Wang et al., 2024h),(Zhang et al., 2024b)
(Hu et al., 2024b),(Cui et al., 2024a), (Xu et al., 2024a),(Liu et al., 2024g)
@@
(Qiao et al., 2024b),(Zhao et al., 2024b)
Vicuna Y (Wang et al., 2024d)
@@
(Wang et al., 2024f)
To investigate the trend of the LLMERS, we visualize the current works based on the type of LLM and semantics in Figure 6. As mentioned before, the key to LLM enhancement lies in the semantics. At the early stage, many works (Xi et al., 2024a; Zhao et al., 2024c) prompt the LLM to derive the natural language, which contains knowledge and semantics for augmentation. Since they are readable and understandable, we denote them as explicit semantics. By comparison, the cluster of implicit semantics means that the mediation from LLM for enhancement is non-verbal, e.g., the hidden states of LLM. Though humans may favor explicit knowledge due to its explainability, we find that recent works prefer the implicit one. The reason lies in the better performance of implicit knowledge while explicit one suffers from information loss during the extra encoding process.
@@
  • Model enhancement has gained more attention recently, because it can combine implicit semantics and fine-tuned LLM.
6. Applications, Efficiency, and Resources
Applications. Traditional recommender systems have been widely adopted in various applications, such as news recommendations. Though LLMERS is also based on traditional RS, it often requires abundant side information, especially informative texts, to make full use of the reasoning and understanding abilities of LLM. Thus, current LLMERS works often focus on applications that have one of two following characteristics: (i) it has a number of normal features that LLM can understand; (ii) it has a volume of texts that LLM can summarize. The E-commerce application belongs to the former one, where users have many profile features and commodities have attributes (Wang et al., 2024a; Liu et al., 2024e). Since the LLM owns open-world knowledge, these features can be understood by it and derive semantics for enhancement. By comparison, the news recommendation is the typical application that utilizes the LLM to summarize the abundant news texts (Liu et al., 2024a). To show it more clearly, we list the typical applications that LLMERS has targeted in Table 2.
Efficiency. Compared with LLM as RS, the most important distinction and merit of LLMERS lie in high efficiency. Thus, we discuss how each category of LLMERS is efficient in the paragraph. i) Knowledge Enhancement minimizes runtime LLM dependency by pre-generating static intermediaries (text summaries or knowledge graphs), but introduces persistent encoding overhead. Summary-based methods append cached summaries as supplementary features via text encoders (Li et al., 2023), while KG follows dual paths: feature-based KG integrates with content-based systems through embeddings (Huang et al., 2024a), and interaction-based KG enhance collaborative filtering via graph processing (Wang et al., 2024a). Hybrid methods combine KG-refined summaries with text encoding, trading compounded costs for semantic enrichment. ii) Interaction Enhancement exclusively amplifies training data through LLM-synthesized user interactions (Yang et al., 2024c), preserving inference efficiency and achieving universal scalability across recommendation architectures. iii) Model Enhancement employs three strategies with varying architectural dependencies: Initialization of components (content-based models (Li et al., 2023) or embedding layers (Hu et al., 2024b)) are enhanced by LLM-generated knowledge, and only small RS models are used for service. Distillation prioritizes traditional RS models (Cui et al., 2024a), eliminating LLM while inference. Embedding Fusion decouples LLMs from inference by caching embeddings as semantic features (content-based RS (Liu et al., 2024b)) or cross-modal alignment anchors (collaborative filtering (Jia et al., 2024)), incurring only storage and view-encoding costs.
Resources. Datasets and benchmarks are often the most vital resources for one research field. For the dataset, many have included either informative features or abundant texts for LLM usage. We list the available public datasets with their access links according to the applications in Table 2 to facilitate further research. As for the benchmark, unfortunately, there has been no one since it is an emerging direction. Nevertheless, many research studies have open-sourced their code, which pushes the advancements of this field. We conclude all the available codes in our survey repository  [^1] https://github.com/liuqidong07/Awesome-LLM-Enhanced-Recommender-Systems.